OpenStudy (anonymous):

Given arithmetic sequences A and B shown, what is the positive difference between the 2011th term of sequence A and the 2011th term of sequence B? Sequence A: 1, 5, 9, 13, . . . Sequence B: 1, 7, 13, 19, . . .

OpenStudy (anonymous):

1 + 2010*6 - (1 + 2010*4) = 2010(6-4) = 4020

OpenStudy (anonymous):

remember terms of an AS are a,a+d,a+2d.....

OpenStudy (anonymous):

nth term = a + (n-1)d

OpenStudy (anonymous):

follow that ok?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

great

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

yw